IGO Night Logs Cycle 2011A

May 2011 (Total available observing time: 140 hours)

Hours Observed: 83 hours
Weather Loss: 41 hours
Technical Loss: 5 hours

Date

Hours Observed

Hours Lost

Comments

Weather

Technical

Others

May 14

4

6

0

0

Cloudy Sky.

May 13

3

7

0

0

Hazy and cloudy sky.

May 12

5

0

1

4

Altitude axis failure. Bright moon.

May 11

9

0

0

1

Observed.

May 10

9

0

0

1

Observed.

May 09

3

7

0

0

Hazy and Cloudy sky.

May 08

3

6

0

1

Hazy and Cloudy sky. Generator problem. Solved afrer some time.

May 07

9

0

0

1

Observed.

May 06

8

0

1

1

Observed. Liquid Nitrogen got over.

May 05

8

0

1

1

Observed. Liquid Nitrogen got over.

May 04

0

10

0

0

Cloudy sky.

May 03

7

0

2

1

Observed. Problem with Cass. Axis and autoguider focus.

May 02

10

0

0

0

Clear sky. Observed.

May 01

5

5

0

0

Sky became clody after midnight.

April 2011 (Total available observing time: 300 hours)

Hours Observed: 168 hours
Weather Loss: 86 hours
Technical Loss: 13 hours

Date

Hours Observed

Hours Lost

Comments

Weather

Technical

Others

Apr 30

0

10

0

0

Cloudy sky.

Apr 29

0

0

10

0

Polarimeter not working.

Apr 28

10

0

0

0

Perhaps polarimeter not working properly.

Apr 27

9

0

1

0

Dome and azimuth problems. Solved during the night.

Apr 26

10

0

0

0

Observed. Thin clouds.

Apr 25

0

10

0

0

Cloudy sky.

Apr 24

0

10

0

0

Clouds, Rain.

Apr 23

4

6

0

0

Cloudy after midnight.

Apr 22

6

0

0

4

Clear. Observed.

Apr 21

6

0

0

4

Clear. Observed.

Apr 20

5

0

0

5

Hazy. Bright moon.

Apr 19

5

0

0

5

Observed. Moonlight.

Apr 18

5

0

0

5

Observed. Moonlight.

Apr 17

10

0

0

0

Observed.

Apr 16

7

3

0

0

Rain at beginning. Clear afterwards. Observed.

Apr 15

10

0

0

0

Observed. Thin clouds.

Apr 14

4

3

0

3

Thin Clouds. Moonlight.

Apr 13

0

10

0

0

Cloudy sky.

Apr 12

0

10

0

0

Cloudy sky.

Apr 11

0

10

0

0

Cloudy sky.

Apr 10

0

10

0

0

Cloudy sky.

Apr 09

7

0

0

3

Observed.

Apr 08

5

4

1

0

Thin Clouds. TCS failed. Cloudy afrer wards.

Apr 07

10

0

0

0

Observed.

Apr 06

10

0

0

0

Observed.

Apr 05

6

0

0

4

Observed.

Apr 04

10

0

0

0

Observed.

Apr 03

9

0

1

0

Observed. AMN failed.

Apr 02

10

0

0

0

Observed. Windy sky. Bad seeing.

Apr 01

10

0

0

0

Observed.



March 2011 (Total available observing time: 310 hours)

Hours Observed: 201 hours
Weather Loss: 32 hours
Technical Loss: 7 hours

Date

Hours Observed

Hours Lost

Comments

Weather

Technical

Others

Mar 31

9

0

0

1

Observed.

Mar 30

10

0

0

0

Observed.

Mar 29

9

0

1

0

Observed. Self oper off. Oper on given again.

Mar 28

10

0

0

0

Observed.

Mar 27

8

0

0

2

Observed.

Mar 26

8

2

0

0

Observed. Sky cloudy afterwards.

Mar 25

0

0

0

10

TIFR Instrument Changeover.

Mar 24

1

9

0

0

Cloudy sky.

Mar 23

4

0

6

0

IT rack and P01 problem. Full system restart given.

Mar 22

7

0

0

3

Observed.

Mar 21

5

0

0

5

Observed. Moonlight.

Mar 20

1

9

0

0

Cloudy sky.

Mar 19

0

0

0

10

TIFR Instrument Changeover.

Mar 18

0

0

0

10

TIFR Instrument Changeover.

Mar 17

0

0

0

10

Pointing Model observations.

Mar 16

0

0

0

10

Pointing Model observations.

Mar 15

10

0

0

0

Observed.

Mar 14

10

0

0

0

Observed.

Mar 13

6

0

0

4

Observed.

Mar 12

7

0

0

3

Observed.

Mar 11

9

0

0

1

Observed.

Mar 10

10

0

0

0

Observed.

Mar 09

9

0

0

1

Observed.

Mar 08

10

0

0

0

Observed.

Mar 07

10

0

0

0

Observed.

Mar 06

8

2

0

0

Observed. Partly Cloudy.

Mar 05

10

0

0

0

Observed.

Mar 04

10

0

0

0

Observed.

Mar 03

10

0

0

0

Observed. Thin clouds.

Mar 02

10

0

0

0

Observed. Thin clouds.

Mar 01

0

10

0

0

Cloudy sky.



February 2011 (Total available observing time: 280 hours)

Hours Observed: 215 hours
Weather Loss: 27 hours
Technical Loss: 01 hours

Date

Hours Observed

Hours Lost

Comments

Weather

Technical

Others

Feb 28

0

10

0

0

Cloudy Sky.

Feb 27

2

8

0

0

Cloudy Sky. Tokk afew observations.

Feb 26

10

0

0

0

Observed.

Feb 25

9

0

1

0

Observed. AMN problem. Oper Off/On had to be given.

Feb 24

6

0

0

4

Observed.

Feb 23

3

7

0

0

Mostly cloudy. Observed for some time.

Feb 22

0

0

0

10

TIFR Instrument Changeover.

Feb 21

10

0

0

0

TIFR Observations.

Feb 20

10

0

0

0

TIFR Observations.

Feb 19

10

0

0

0

TIFR Observations.

Feb 18

10

0

0

0

TIFR Observations.

Feb 17

9

1

0

0

Thin Clouds. Observed.

Feb 16

9

1

0

0

Initially Cloudy. Observed.

Feb 15

0

0

0

10

TIFR Instrument Changeover.

Feb 14

0

0

0

10

TIFR Instrument Changeover.

Feb 13

10

0

0

0

Observed.

Feb 12

7

0

0

3

Observed.

Feb 11

10

0

0

0

Observed.

Feb 10

10

0

0

0

Observed.

Feb 09

10

0

0

0

Observed.

Feb 08

10

0

0

0

Observed.

Feb 07

10

0

0

0

Observed.

Feb 06

10

0

0

0

Observed.

Feb 05

10

0

0

0

Observed.

Feb 04

10

0

0

0

Observed.

Feb 03

10

0

0

0

Observed.

Feb 02

10

0

0

0

Observed.

Feb 01

10

0

0

0

Observed.